1.)  How is the unemployment rate measured? Are part-time workers who want to work full-time considered officially “unemployed” in the rate?

2.) Consumer spending is one component of the economy as calculated by GDP. What are the others? How is economic growth measured?

3.) “Over the past year, wage growth ticked up just 2.9%, which was lower than expected and the weakest annual growth since July 2018." Over the same time period, inflation was 1.7%. As a result, what was the “real” wage growth in the economy over the past year? How did you calculate it and what does it mean?

4.) Following the jobs report on Friday, Fed Chairman Jerome Powell said the economy is in "a good place. Our job is to keep it there," he added.” Do you think the U.S. economy will continue to grow? Why or why not? What examples can you provide to support your answer?
Answer #1

1. The unemployment rate is calculated by the formula

Unemployment rate = number of people unemployed ÷ labour force

It is expressed in percentage. So the obtained value from the above formula is multiplied by 100.

Labour force is the total of both employed and unemployed people.

The part time workers who are willing to work full time can be considered as involuntary part time workers. Thus they are not considered as officially unemployed.

2. The other components of GDP are government spending, investment and difference between the exports and imports.

Economic growth is measured by the formula

GDP = C+ I + G + ( X_M )

Where ,

GDP = economic growth

C = consumption

I = investment

G = government spending

(X_M) = exports _ imports.

Consumption means the expenditure by consumers to meet their needs.

Investment comprises of goods that we consume in the future

Government spending are the expenses incurred by the government for improving the standard of living of the people.

3. The real wage growth in the economy is calculated using the formula

Real Wage = wage rate ÷ inflation

Given,

Wage rate = 2.9 %

Inflation = 1.7%

Real Wage = 2.9 ÷ 1.7 = 1.7%

The sum total of goods and services that can be purchased with the wages is known as real wage. It is inflation adjusted rate.

4. The economy of the US will grow but that growth will be at a decreasing rate.

The factors that lowers growth rate are

  • Trade war with China aiming to reduce the trade deficit with China. This will negatively impact both the US and China's economy
  • Rising unemployment rate.
  • Rising prices, inflation

All these factors pull down the increase in growth rate